# From Perception to Reasoning: Deep Thinking Empowers Multimodal Large Language Models

### v1

@@ 5.2 Autonomous Driving
In practical system construction, Dolphins(ma2024dolphins), a conversational driving assistant system, introduces a M-CoT process based on scene context to enhance its understanding and execution of driving instructions. This system is fine-tuned on datasets specifically designed for intelligent driving, significantly improving its performance in intelligent cockpits and human–machine interaction. Agent-Driver(mao2023language) employs a large language model as the core cognitive agent and transforms traditional autonomous driving pipelines by integrating general-purpose toolkits, achieving end-to-end modeling of the task chain from perception and understanding to decision-making.
Regarding data resources and model validation, several key studies have developed open datasets that support multimodal reasoning and end-to-end driving tasks. DriveCoT(wang2024drivecot) trained the baseline DriveCoT-Agent on its constructed dataset, marking the first successful integration of multimodal CoT methods into an end-to-end autonomous driving system. Reason2Drive (nie2024reason2drive) provides over 600,000 video-text pairs containing explicit reasoning processes, greatly advancing research on interpretable reasoning in autonomous driving. DriveLMM-o1(ishaq2025drivelmm) focuses on stepwise visual reasoning tasks, covering three critical aspects in autonomous driving scenes: perception, prediction, and planning, thus offering datasets and benchmarks for systematic evaluation of multimodal reasoning models.
In terms of reasoning frameworks, PKRD-CoT(luo2024pkrd) proposes a unified prompt design framework for multimodal large language models in autonomous driving, guiding models through a human-like stepwise thinking process to make real-time driving decisions, thereby enhancing system adaptability and responsiveness. OpenEMMA(xing2025openemma) integrates CoT reasoning mechanisms with external expert vision models, demonstrating outstanding performance in vehicle trajectory planning tasks, particularly in 3D structural recognition and precise detection of road targets. Meanwhile, the study by Islam et al.(islam2024application) decomposes the autonomous driving decision-making process into three stages: scene understanding, behavior prediction, and decision generation, introducing CoT as the core pathway linking reasoning across these stages, thereby improving the overall decision-making rationality and coherence.
@@ 5.3 Healthcare
In medical image and video analysis, Chain-of-Look(xi2023chain) models the triplet recognition task in endoscopic videos as a visual prompt generation problem by leveraging vision-language models. It explicitly decomposes the task into a sequence of temporally structured video reasoning steps. This approach not only improves diagnostic accuracy but also strengthens the model’s ability to localize key pathological features within image sequences.
For medical visual question answering (VQA) tasks, MedThink(gai2024medthink) proposes simultaneously generating decision outcomes and corresponding reasoning rationales, intuitively demonstrating the system’s reasoning pathways in answering medical questions, thereby enhancing the interpretability of the decision process. MedCoT(liu2024medcot) introduces a hierarchical expert verification CoT approach that simulates multidisciplinary clinical consultation by integrating multiple “expert models” for reasoning and correction, achieving more reliable and clinically consistent answers. Additionally, it employs a multi-expert mixed review mechanism to effectively mitigate uncertainties caused by ambiguity or incomplete information in medical Q&A.
In the mental health domain, Dai et al.(dai2024interpretable) apply CoT reasoning methods to stress detection systems, mimicking the cognitive and judgment processes of psychological experts. This enables the model to attain higher interpretability and user trust when assessing highly subjective psychological states.
Regarding medical procedural action recognition, TI-PREGO(plini2024ti) combines in-context learning (ICL) with an automatic chain-of-thought (ACoT) mechanism to perform procedural error detection in egocentric medical videos through action recognition and prediction tasks.
Meanwhile, MedVLM-R1(pan2025medvlm) enhances the reasoning ability of vision-language models in medical image analysis tasks by incorporating reinforcement learning, explicitly strengthening the transparency and traceability of the reasoning paths through natural language reasoning.
@@ 5.4 Multimodal Generation
In terms of system architecture, the MINT model(wang2025mint) introduces a parallel ensemble structure of multiple expert modules—MTXpert (Mixture of Transformer Experts)—which efficiently integrates natural language generation capabilities with image understanding, avoiding generation degradation caused by modal conflicts. MINT adopts a multi-stage CoT paradigm of “thinking—reasoning—reflecting” during generation, substantially improving the semantic consistency and visual quality of generated images. Regarding physical consistency modeling, the Phys-AR framework(lin2025reasoning) combines symbolic reasoning with reinforcement learning mechanisms to construct a multimodal generation system endowed with causal understanding capabilities.
Furthermore, in the context of generation authenticity and traceability, the CoT-Finetuned model (agrahari2025tracing) explores the application of chain-of-thought reasoning mechanisms to AI-generated content detection tasks. This model not only determines whether text is generated by large language models but also identifies the specific model used, thereby effectively enhancing the interpretability and discriminative power of detection systems. This approach holds significant importance for maintaining academic integrity, combating misinformation, and advancing AI ethics standards.
@@ 5.5 Machine Translation
CoT-ST(du2024cot) leverages multimodal chain-of-thought by decomposing speech translation into sequential steps of speech recognition and translation, enhancing the natural synergy and complementarity between these subtasks. This leads to significant improvements in translation performance of speech language models (SLMs). IMAGE(chen2024make) integrates multimodal large models with stable diffusion(rombach2022high)-based networks to explicitly generate images for each sentence to be translated, utilizing visual information to boost machine translation performance. Furthermore, in multimodal translation research targeting low-resource languages, Rajpoot et al.(rajpoot2024multimodal) propose a translation strategy that combines image captioning with chain-of-thought reasoning. This method extracts semantic descriptions from images through multiple stages and embeds them into the reasoning path of the language translation process, achieving promising results in English-to-low-resource Indian language multimodal translation tasks.
@@ 5.6 Affective Computing and Social Reasoning
In affective computing, MM-PEAR-CoT(li2025multimodal) is the first to introduce the chain-of-thought reasoning paradigm into multimodal emotion recognition tasks. By fusing textual, visual, and audio information, it alleviates reasoning biases caused by unimodal errors in traditional chain-of-thought methods, thereby achieving more stable and interpretable performance in emotion label classification tasks. Additionally, the Empatheia system(zhang2025towards) integrates multimodal large language models with empathetic reasoning chains, demonstrating more natural, nuanced, and personalized language generation capabilities in empathetic response generation (e.g., mental health Q&A, emotional counseling), thus advancing the practical application of intelligent systems in social-affective interaction.
Regarding social reasoning, SOCIAL GENOME(mathur2025social) offers the first benchmark framework to evaluate multimodal models’ social reasoning capabilities, featuring fine-grained reasoning steps supplemented with external knowledge concepts for auxiliary assessment. Yu and Luo(yu2024chain) further explore the application of multimodal large language models enhanced with chain-of-thought prompting methods in demographic inference, significantly reducing prediction bias. Moreover, Hu et al.(hu2024multimodal) focus on child physical and mental health protection by applying multimodal chain-of-thought for age-appropriate rating of mobile applications. CPFEND(xu2024multimodal) and RAGAR(khaliq2024ragar) are devoted to improving models’ reasoning and judgment capabilities in misinformation detection and political fact-checking.
@@ 6.1 Robust Reasoning
The hallucination problem in large language models has long troubled researchers. Traditional large language models, due to the absence of explicit reasoning processes, frequently produce contextually inconsistent or factually inaccurate responses, severely limiting their real-world applications. Although chain-of-thought reasoning alleviates hallucinations to some extent, it also introduces side effects such as “overthinking.” Common hallucination mitigation strategies during the reasoning phase include external knowledge-based methods (e.g., retrieval-augmented generation, expert model-assisted detection(zhang2024meter; zhang2025critic)) and internal mechanism-based methods (e.g., self-reflection, self-consistency reasoning). However, these methods mostly operate at the overall answer level, suffering from error accumulation across intermediate reasoning steps. Future research could explore finer-grained optimization at the step level; for instance, HaluSearch(cheng2025think) models answer generation as an explicit stepwise reasoning process using Monte Carlo tree search and leverages step-level reward mechanisms to explore optimal reasoning paths, thereby enhancing reasoning reliability.
@@
Moreover, in multimodal environments, insufficient alignment between different modalities can also trigger hallucinations. For example, Zheng et al.(zheng2024thinking) pointed out that the commonly used “look-then-think” paradigm in current multimodal chain-of-thought methods—where reasoning chains are generated simultaneously with visual input—can be misled by deceptive visual information, resulting in biased reasoning outcomes. Therefore, cross-modal alignment mechanisms in future multimodal chain-of-thought research will become a key focus. For instance, the VIC framework(zheng2024thinking) constructs the reasoning chain based on textual context prior to introducing visual information, thereby reducing cross-modal bias caused by visual interference and improving reasoning accuracy.
@@
Apart from hallucinations, the length of the reasoning chain also directly impacts reasoning robustness. A chain that is too short may lead to insufficient deliberation, affecting the accuracy of the final conclusion; conversely, an excessively long chain can cause information forgetting and overthinking. Hence, designing effective mechanisms for controlling reasoning chain length is a promising research direction. Some existing works have proposed solutions in this regard. For example, Sun et al.(sun2025mitigating) addressed the issue of gradually diminished attention to visual information during the reasoning process in multimodal large language models by proposing the TVC (Take-along Visual Conditioning) strategy, which maintains image information into critical reasoning stages and employs dynamic pruning to compress redundant visual representations, thereby sustaining the model’s focus on visual content. Meanwhile, ThinkEdit(sun2025thinkedit) tackles performance degradation caused by overly short reasoning chains in mathematical problem solving by introducing a weight editing method to mitigate this issue. Yang et al.(yang2025towards) observed that excessively long chains weaken mathematical reasoning capabilities and thus proposed the TOPS (Thinking-Optimal Scaling) strategy, enabling the model to self-optimize and arrive at correct answers with shorter reasoning chains. Furthermore, Yang et al.(yang2025dynamic) proposed a plug-and-play mechanism that detects the model’s behavior at potential reasoning turning points (e.g., “wait” tokens) and dynamically terminates subsequent chain generation when the model shows high confidence in tentative answers, thus enabling more efficient and rational planning of reasoning paths.
@@ 6.2 Secure Reasoning
The safety and ethical issues of reasoning models are critical topics that urgently need to be addressed. Due to their characteristics of cross-modal alignment and explicit step-by-step reasoning, multimodal reasoning models exhibit unique vulnerabilities, making them more susceptible to malicious attacks and the generation of harmful information. Zhou et al.(zhou2025hidden) evaluated the robustness of mainstream large reasoning models against adversarial attacks such as jailbreaking and prompt injection, finding that open-source models like DeepSeek-R1 and distilled models have security disadvantages. In multimodal scenarios, Fang et al.(fang2025safemlrm) revealed the ”Reasoning Tax” phenomenon, wherein a model’s acquisition of complex reasoning capabilities can weaken the safety alignment features of the base model. At the same time, the study found that multimodal reasoning models possess a certain capacity for self-correction when faced with unsafe reasoning chains.
@@
Currently, research on specific attack methods and defense mechanisms for multimodal reasoning models is still in its nascent stages, whereas related research on unimodal (text-only) reasoning models or general-purpose multimodal models is relatively more mature. Wang et al.(wang2025safety) categorized the attack methods for large reasoning models into four types: reasoning length attacks, answer correctness attacks, prompt injection attacks, and jailbreak attacks. Reasoning length attacks induce the model to engage in ”overthinking”(kumar2025overthink; zaremba2025trading) or ”underthinking”(zaremba2025trading) through prompts, where overthinking attacks are essentially a form of Denial of Service (DoS) attack. Answer correctness attacks tamper with intermediate reasoning steps through backdoor attacks(xiang2024badchain; guo2025darkmind; zhao2025shadowcot) or error injection(cui2025process) to compromise the integrity of the results. Prompt injection attacks disguise malicious instructions as normal user input to bypass the model’s preset safety protocols(kumar2024strengthening; liu2023prompt). Lastly, jailbreak attacks use prompts to evade the model’s safety restrictions. Some studies(russinovich2025great; ren2025llms; li2024llm) also utilize the multi-turn interaction of CoT to progressively guide the model to generate harmful content.
@@
For vision-language models, Ye et al. classified attack methods into three categories based on the attacker’s knowledge of the model architecture: white-box, gray-box, and black-box(ye2025survey). White-box attacks(qi2024visual; bailey2023image; wang2024white) occur when the attacker has full knowledge of the model’s internal architecture, amplifying the attack’s effect by injecting adversarial noise into images and modifying text inputs to exploit modal dependencies. Gray-box attacks(zhao2023evaluating; wang2024break) leverage partial knowledge of the model, using open-source vision encoders like CLIP to generate adversarial samples. Black-box attacks(gong2025figstep; ma2024visual) bypass safety restrictions by designing clever prompts without any knowledge of the model’s internal information.
@@
In terms of defense mechanisms, current research primarily focuses on three levels: training-time defense, inference-time defense, and external guardrail models(wang2025safety). During the model alignment training phase, some work is dedicated to constructing Chain-of-Thought safety datasets(jiang2025safechain; zhang2025realsafe; chen2024dress) and performing safety reasoning alignment through supervised fine-tuning(lou2025think; xia2025msr) or reinforcement learning(guan2024deliberative; zhang2025stair). To avoid the safety alignment tax—where safety training may sacrifice model performance—researchers are exploring the application of techniques at the inference stage, such as safety prompts(wang2024adashield; zhao2024bluesuffix), the detection and removal of adversarial noise(sun2024safeguarding), and safe decoding(jiang2025safechain), to enhance model security. The guardrail model approach(liu2025guardreasoner; wen2025thinkguard) utilizes an independent auditing model to monitor and filter the inputs and outputs of the main model in real-time, thereby achieving security protection without modifying the original model.

### v2

@@ 5.2 Autonomous Driving
In practical system construction, Dolphins(Ma et al., 2024b), a conversational driving assistant system, introduces a M-CoT process based on scene context to enhance its understanding and execution of driving instructions. This system is fine-tuned on datasets specifically designed for intelligent driving, significantly improving its performance in intelligent cockpits and human–machine interaction. Agent-Driver(Mao et al., 2023) employs a large language model as the core cognitive agent and transforms traditional autonomous driving pipelines by integrating general-purpose toolkits, achieving end-to-end modeling of the task chain from perception and understanding to decision-making.
Regarding data resources and model validation, several key studies have developed open datasets that support multimodal reasoning and end-to-end driving tasks. DriveCoT(Wang et al., 2024d) trained the baseline DriveCoT-Agent on its constructed dataset, marking the first successful integration of multimodal CoT methods into an end-to-end autonomous driving system. Reason2Drive (Nie et al., 2024) provides over 600,000 video-text pairs containing explicit reasoning processes, greatly advancing research on interpretable reasoning in autonomous driving. DriveLMM-o1(Ishaq et al., 2025) focuses on stepwise visual reasoning tasks, covering three critical aspects in autonomous driving scenes: perception, prediction, and planning, thus offering datasets and benchmarks for systematic evaluation of multimodal reasoning models.
In terms of reasoning frameworks, PKRD-CoT(Luo et al., 2024) proposes a unified prompt design framework for multimodal large language models in autonomous driving, guiding models through a human-like stepwise thinking process to make real-time driving decisions, thereby enhancing system adaptability and responsiveness. OpenEMMA(Xing et al., 2025) integrates CoT reasoning mechanisms with external expert vision models, demonstrating outstanding performance in vehicle trajectory planning tasks, particularly in 3D structural recognition and precise detection of road targets. Meanwhile, the study by Islam et al.(Islam, 2024) decomposes the autonomous driving decision-making process into three stages: scene understanding, behavior prediction, and decision generation, introducing CoT as the core pathway linking reasoning across these stages, thereby improving the overall decision-making rationality and coherence.
@@ 5.3 Healthcare
In medical image and video analysis, Chain-of-Look(Xi et al., 2023) models the triplet recognition task in endoscopic videos as a visual prompt generation problem by leveraging vision-language models. It explicitly decomposes the task into a sequence of temporally structured video reasoning steps. This approach not only improves diagnostic accuracy but also strengthens the model’s ability to localize key pathological features within image sequences.
For medical visual question answering (VQA) tasks, MedThink(Gai et al., 2024) proposes simultaneously generating decision outcomes and corresponding reasoning rationales, intuitively demonstrating the system’s reasoning pathways in answering medical questions, thereby enhancing the interpretability of the decision process. MedCoT(Liu et al., 2024a) introduces a hierarchical expert verification CoT approach that simulates multidisciplinary clinical consultation by integrating multiple “expert models” for reasoning and correction, achieving more reliable and clinically consistent answers. Additionally, it employs a multi-expert mixed review mechanism to effectively mitigate uncertainties caused by ambiguity or incomplete information in medical Q&A.
In the mental health domain, Dai et al.(Dai, 2024) apply CoT reasoning methods to stress detection systems, mimicking the cognitive and judgment processes of psychological experts. This enables the model to attain higher interpretability and user trust when assessing highly subjective psychological states.
Regarding medical procedural action recognition, TI-PREGO(Plini et al., 2024) combines in-context learning (ICL) with an automatic chain-of-thought (ACoT) mechanism to perform procedural error detection in egocentric medical videos through action recognition and prediction tasks.
Meanwhile, MedVLM-R1(Pan et al., 2025) enhances the reasoning ability of vision-language models in medical image analysis tasks by incorporating reinforcement learning, explicitly strengthening the transparency and traceability of the reasoning paths through natural language reasoning.
@@ 5.4 Multimodal Generation
In terms of system architecture, the MINT model(Wang et al., 2025g) introduces a parallel ensemble structure of multiple expert modules—MTXpert (Mixture of Transformer Experts)—which efficiently integrates natural language generation capabilities with image understanding, avoiding generation degradation caused by modal conflicts. MINT adopts a multi-stage CoT paradigm of “thinking—reasoning—reflecting” during generation, substantially improving the semantic consistency and visual quality of generated images. Regarding physical consistency modeling, the Phys-AR framework(Lin et al., 2025b) combines symbolic reasoning with reinforcement learning mechanisms to construct a multimodal generation system endowed with causal understanding capabilities.
Furthermore, in the context of generation authenticity and traceability, the CoT-Finetuned model (Agrahari & Singh, 2025) explores the application of chain-of-thought reasoning mechanisms to AI-generated content detection tasks. This model not only determines whether text is generated by large language models but also identifies the specific model used, thereby effectively enhancing the interpretability and discriminative power of detection systems. This approach holds significant importance for maintaining academic integrity, combating misinformation, and advancing AI ethics standards.
@@ 5.5 Machine Translation
CoT-ST(Du et al., 2024) leverages multimodal chain-of-thought by decomposing speech translation into sequential steps of speech recognition and translation, enhancing the natural synergy and complementarity between these subtasks. This leads to significant improvements in translation performance of speech language models (SLMs). IMAGE(Chen et al., 2024a) integrates multimodal large models with stable diffusion(Rombach et al., 2022)-based networks to explicitly generate images for each sentence to be translated, utilizing visual information to boost machine translation performance. Furthermore, in multimodal translation research targeting low-resource languages, Rajpoot et al.(Rajpoot et al., 2024) propose a translation strategy that combines image captioning with chain-of-thought reasoning. This method extracts semantic descriptions from images through multiple stages and embeds them into the reasoning path of the language translation process, achieving promising results in English-to-low-resource Indian language multimodal translation tasks.
@@ 5.6 Affective Computing and Social Reasoning
In affective computing, MM-PEAR-CoT(Li et al., 2025b) is the first to introduce the chain-of-thought reasoning paradigm into multimodal emotion recognition tasks. By fusing textual, visual, and audio information, it alleviates reasoning biases caused by unimodal errors in traditional chain-of-thought methods, thereby achieving more stable and interpretable performance in emotion label classification tasks. Additionally, the Empatheia system(Zhang et al., 2025d) integrates multimodal large language models with empathetic reasoning chains, demonstrating more natural, nuanced, and personalized language generation capabilities in empathetic response generation (e.g., mental health Q&A, emotional counseling), thus advancing the practical application of intelligent systems in social-affective interaction.
Regarding social reasoning, SOCIAL GENOME(Mathur et al., 2025) offers the first benchmark framework to evaluate multimodal models’ social reasoning capabilities, featuring fine-grained reasoning steps supplemented with external knowledge concepts for auxiliary assessment. Yu and Luo(Yu & Luo, 2024) further explore the application of multimodal large language models enhanced with chain-of-thought prompting methods in demographic inference, significantly reducing prediction bias. Moreover, Hu et al.(Hu et al., 2024a) focus on child physical and mental health protection by applying multimodal chain-of-thought for age-appropriate rating of mobile applications. CPFEND(Xu et al., 2024b) and RAGAR(Khaliq et al., 2024) are devoted to improving models’ reasoning and judgment capabilities in misinformation detection and political fact-checking.
@@ 6.1 Robust Reasoning
The hallucination problem in large language models has long troubled researchers. Traditional large language models, due to the absence of explicit reasoning processes, frequently produce contextually inconsistent or factually inaccurate responses, severely limiting their real-world applications. Although chain-of-thought reasoning alleviates hallucinations to some extent, it also introduces side effects such as “overthinking.” Common hallucination mitigation strategies during the reasoning phase include external knowledge-based methods (e.g., retrieval-augmented generation, expert model-assisted detection(Zhang et al., 2024f; 2025c)) and internal mechanism-based methods (e.g., self-reflection, self-consistency reasoning). However, these methods mostly operate at the overall answer level, suffering from error accumulation across intermediate reasoning steps. Future research could explore finer-grained optimization at the step level; for instance, HaluSearch(Cheng et al., 2025a) models answer generation as an explicit stepwise reasoning process using Monte Carlo tree search and leverages step-level reward mechanisms to explore optimal reasoning paths, thereby enhancing reasoning reliability.
@@
Moreover, in multimodal environments, insufficient alignment between different modalities can also trigger hallucinations. For example, Zheng et al.(Zheng et al., 2024b) pointed out that the commonly used “look-then-think” paradigm in current multimodal chain-of-thought methods—where reasoning chains are generated simultaneously with visual input—can be misled by deceptive visual information, resulting in biased reasoning outcomes. Therefore, cross-modal alignment mechanisms in future multimodal chain-of-thought research will become a key focus. For instance, the VIC framework(Zheng et al., 2024b) constructs the reasoning chain based on textual context prior to introducing visual information, thereby reducing cross-modal bias caused by visual interference and improving reasoning accuracy.
@@
Apart from hallucinations, the length of the reasoning chain also directly impacts reasoning robustness. A chain that is too short may lead to insufficient deliberation, affecting the accuracy of the final conclusion; conversely, an excessively long chain can cause information forgetting and overthinking. Hence, designing effective mechanisms for controlling reasoning chain length is a promising research direction. Some existing works have proposed solutions in this regard. For example, Sun et al.(Sun et al., 2025b) addressed the issue of gradually diminished attention to visual information during the reasoning process in multimodal large language models by proposing the TVC (Take-along Visual Conditioning) strategy, which maintains image information into critical reasoning stages and employs dynamic pruning to compress redundant visual representations, thereby sustaining the model’s focus on visual content. Meanwhile, ThinkEdit(Sun et al., 2025a) tackles performance degradation caused by overly short reasoning chains in mathematical problem solving by introducing a weight editing method to mitigate this issue. Yang et al.(Yang et al., 2025b) observed that excessively long chains weaken mathematical reasoning capabilities and thus proposed the TOPS (Thinking-Optimal Scaling) strategy, enabling the model to self-optimize and arrive at correct answers with shorter reasoning chains. Furthermore, Yang et al.(Yang et al., 2025a) proposed a plug-and-play mechanism that detects the model’s behavior at potential reasoning turning points (e.g., “wait” tokens) and dynamically terminates subsequent chain generation when the model shows high confidence in tentative answers, thus enabling more efficient and rational planning of reasoning paths.
@@ 6.2 Secure Reasoning
The safety and ethical issues of reasoning models are critical topics that urgently need to be addressed. Due to their characteristics of cross-modal alignment and explicit step-by-step reasoning, multimodal reasoning models exhibit unique vulnerabilities, making them more susceptible to malicious attacks and the generation of harmful information. Zhou et al.(Zhou et al., 2025) evaluated the robustness of mainstream large reasoning models against adversarial attacks such as jailbreaking and prompt injection, finding that open-source models like DeepSeek-R1 and distilled models have security disadvantages. In multimodal scenarios, Fang et al.(Fang et al., 2025) revealed the ”Reasoning Tax” phenomenon, wherein a model’s acquisition of complex reasoning capabilities can weaken the safety alignment features of the base model. At the same time, the study found that multimodal reasoning models possess a certain capacity for self-correction when faced with unsafe reasoning chains.
@@
Currently, research on specific attack methods and defense mechanisms for multimodal reasoning models is still in its nascent stages, whereas related research on unimodal (text-only) reasoning models or general-purpose multimodal models is relatively more mature. Wang et al.(Wang et al., 2025a) categorized the attack methods for large reasoning models into four types: reasoning length attacks, answer correctness attacks, prompt injection attacks, and jailbreak attacks. Reasoning length attacks induce the model to engage in ”overthinking”(Kumar et al., 2025; Zaremba et al., 2025) or ”underthinking”(Zaremba et al., 2025) through prompts, where overthinking attacks are essentially a form of Denial of Service (DoS) attack. Answer correctness attacks tamper with intermediate reasoning steps through backdoor attacks(Xiang et al., 2024b; Guo & Tourani, 2025; Zhao et al., 2025a) or error injection(Cui et al., 2025a) to compromise the integrity of the results. Prompt injection attacks disguise malicious instructions as normal user input to bypass the model’s preset safety protocols(Kumar et al., 2024; Liu et al., 2023c). Lastly, jailbreak attacks use prompts to evade the model’s safety restrictions. Some studies(Russinovich et al., 2025; Ren et al., 2025; Li et al., 2024a) also utilize the multi-turn interaction of CoT to progressively guide the model to generate harmful content.
@@
For vision-language models, Ye et al. classified attack methods into three categories based on the attacker’s knowledge of the model architecture: white-box, gray-box, and black-box(Ye et al., 2025). White-box attacks(Qi et al., 2024; Bailey et al., 2023; Wang et al., 2024c) occur when the attacker has full knowledge of the model’s internal architecture, amplifying the attack’s effect by injecting adversarial noise into images and modifying text inputs to exploit modal dependencies. Gray-box attacks(Zhao et al., 2023; Wang et al., 2024i) leverage partial knowledge of the model, using open-source vision encoders like CLIP to generate adversarial samples. Black-box attacks(Gong et al., 2025; Ma et al., 2024a) bypass safety restrictions by designing clever prompts without any knowledge of the model’s internal information.
@@
In terms of defense mechanisms, current research primarily focuses on three levels: training-time defense, inference-time defense, and external guardrail models(Wang et al., 2025a). During the model alignment training phase, some work is dedicated to constructing Chain-of-Thought safety datasets(Jiang et al., 2025b; Zhang et al., 2025g; Chen et al., 2024e) and performing safety reasoning alignment through supervised fine-tuning(Lou et al., 2025; Xia et al., 2025a) or reinforcement learning(Guan et al., 2024; Zhang et al., 2025h). To avoid the safety alignment tax—where safety training may sacrifice model performance—researchers are exploring the application of techniques at the inference stage, such as safety prompts(Wang et al., 2024h; Zhao et al., 2024b), the detection and removal of adversarial noise(Sun et al., 2024b), and safe decoding(Jiang et al., 2025b), to enhance model security. The guardrail model approach(Liu et al., 2025a; Wen et al., 2025) utilizes an independent auditing model to monitor and filter the inputs and outputs of the main model in real-time, thereby achieving security protection without modifying the original model.